1. Consider the case of The Electric Company which produces electricity in New York State. The average monthly demand curve for the firm can be represented by P=65-Q where Q represents the quantity of electricity produced, in megawatt-hours (mwh) and P is measured in cents. Their marginal costs can be represented by MC=5+0.5*Q. Please provide graphs to accompany your analysis.

  1. c. Suppose instead that the firm’s marginal cost curve is more complicated. The firm has two plants. The first plant has a constant marginal cost of 5 cents and can produce up to 20 mwh.  The second plant has a constant marginal cost of 15 cents and can produce up to 10 mwh. Suppose also that the firm faces strongly seasonal demand such that in the summer months demand can be represented by P=70-Q and in the winter months demand can be represented by P=40-Q.  How much will they produce in each season and what prices will they charge? [HINT: Carefully graphing this situation will be VERY HELPFUL]
Answer #1

Answer: a) Given, P=65-Q TR (Total Revenue) = P * Q (65-Q) * Q MR (Marginal Revenue-.NTR / ΔQ => MR 65-2O MC = 5 0.50 [Given]Prite MC 41 21 electri屿(mt)